### v1

# Hybrid Training for Enhanced Multi-task Generalization in Multi-agent Reinforcement Learning

Abstract
In multi-agent reinforcement learning (MARL), achieving multi-task generalization to diverse agents and objectives presents significant challenges. Existing online MARL algorithms primarily focus on single-task performance, but their lack of multi-task generalization capabilities typically results in substantial computational waste and limited real-life applicability. Meanwhile, existing offline multi-task MARL approaches are heavily dependent on data quality, often resulting in poor performance on unseen tasks. In this paper, we introduce HyGen, a novel hybrid MARL framework, Hybrid Training for Enhanced Multi-Task Generalization, which integrates online and offline learning to ensure both multi-task generalization and training efficiency. Specifically, our framework extracts potential general skills from offline multi-task datasets. We then train policies to select the optimal skills under the centralized training and decentralized execution paradigm (CTDE). During this stage, we utilize a replay buffer that integrates both offline data and online interactions. We empirically demonstrate that our framework effectively extracts and refines general skills, yielding impressive generalization to unseen tasks. Comparative analyses on the StarCraft multi-agent challenge show that HyGen outperforms a wide range of existing solely online and offline methods.
Introduction
Figure 1: The overall framework of HyGen is structured as follows: (1) Initially, HyGen learns a global trajectory encoder and action decoders from multi-task data to discover general skills applicable across different tasks. (2) HyGen then learns high-level policies utilizing a hybrid replay buffer that incorporates both offline data and online interactions, essentially refining the skills discovered in the initial stage. (3) During zero-shot execution, HyGen selects and sequences these skills based on a high-level policy and decodes specific actions through the action decoder.
Multi-agent reinforcement learning (MARL) has drawn broad attention for addressing problems in areas such as multi-robot systems [22, 53], video game AIs [35, 4], and autonomous driving [59]. Most existing MARL algorithms remain narrow, in that they focus on optimizing performance for specific tasks [30, 46], resulting in a significant gap between their poor multi-task generalization abilities and the variability of MARL tasks in real-world scenarios. Training specific agents from scratch for each task using MARL algorithms remains very costly and inefficient. Therefore, developing a generalized multi-task MARL algorithm is crucial to address these inefficiencies and improve scalability across diverse MARL tasks.
Two significant obstacles currently limit generalization in multi-task MARL. First, the restrictive model architectures in most MARL algorithms, characterized by fixed input and output dimensions of their neural architectures, fail to accommodate the variability of inputs and outputs across different tasks [20]. Recent online multi-task MARL works primarily focus on training across a predefined set of tasks simultaneously [34, 23] or on fine-tuning pre-trained policies for specific target tasks [20, 63, 36]. Although these approaches utilize a universal input network architecture to address the first obstacle and show promising performance on certain tasks, they fail to resolve another issue of significantly varying policies across different tasks. This results in their learned policies being limited to training tasks and unable to transfer knowledge from source to unseen tasks without further fine-tuning. Offline multi-task MARL [60] involves extracting skills from static datasets and training policies that select and reuse these skills in new tasks, underscoring the potential of leveraging generalizable skills from offline data. However, the effectiveness of these offline methods is often sensitive to the quality of their training dataset. Specifically, when the dataset lacks sufficient optimal or diverse trajectories, agents struggle to learn general skills and optimal policies for source tasks, limiting their performance and generalization capabilities in new tasks. Recent advancements in hybrid reinforcement learning (RL) [12, 19, 50] have shown that extracting skills or behaviors from offline data and then reusing them in online single-agent environments offers a potential solution to addressing issues existing in currently purely online or offline multi-task MARL approaches. However, applications of such frameworks in multi-agent systems remain rare.
In this paper, we propose HyGen: Hybrid Training for Enhanced Multi-Task Generalization, a novel hybrid multi-task MARL approach combining both online environment interaction and offline datasets, as depicted in Figure 1. HyGen first extracts general skills from multi-task offline datasets, and then relies on hybrid training to learn generalizable policies for selecting optimal skills. These general skills and trained policies can then be applied to unseen tasks, as depicted in Figure 1. Specifically, HyGen first extracts general skills using a global trajectory encoder and actions encoders. The global trajectory encoder extracts a set of general skills common across different tasks from multi-task offline datasets, while the action decoders learn to delineate different agent actions with the discovered skills. We then train policies to select the optimal skills to maximize the global return via the centralized training and decentralized execution paradigm (CTDE). During this stage, we utilize a replay buffer that integrates both offline data and online interaction experiences, refining the action decoders to make our skills unconstrained by prior data. Our proposed hybrid training paradigm is unique in how it integrates online interactions with offline data. Unlike RLPD [2], which uses a fixed ratio to sample data from both the online replay buffer and the offline data buffer, our method employs a linearly decreasing ratio. This strategy leverages the efficiency of offline learning initially and progressively incorporates the diversity of online interactions for exploration. Our refined skills during hybrid training are significant compared to previous works in offline multi-task multi-agent skill discovery [60], where skills are constrained to offline data. We finally present empirical results on the Starcraft Multi-Agent Challenge (SMAC), where we show that HyGen achieves remarkable generalization to unseen tasks by discovering general skills and learning high-quality policies, outperforming existing state-of-the-art online and offline multi-task MARL methods.
Related Works
Multi-task MARL
Multi-task MARL methods are more adaptable and efficient than single-task MARL due to knowledge reuse [44] across various tasks. However, reusing knowledge across different tasks comes with its own set of challenges, e.g., varying input and output dimensions, which requires networks with flexible structures, such as self-attention mechanisms [20, 63, 60]. The MT-MARL approach described in [34] distills single-task policies into a unified policy that excels across multiple related tasks. REFIL [23] employs randomized entity-wise factorization for multi-task learning. However, these online methods require simultaneous training across a predefined set of tasks, incurring high costs of online interactions. UPDeT [20] leverages transformer-based value networks to accommodate changes in populations and inputs but requires additional online fine-tuning for new tasks. ODIS [60], an offline multi-task skill-based MARL method, also utilizes transformer-based networks. While this approach generalizes to unseen tasks without additional fine-tuning by reusing general skills, its performance is often constrained by the quality of the dataset. Achieving great generalization ability in unseen tasks remains a challenge.
Skill Discovery in MARL
Skill discovery is an effective approach for tackling complex tasks due to its ability to identify and build a library of skills, often without relying on extrinsic rewards [10]. Recently, single-agent skill learning methods have been extended to MARL. Most skill-based MARL approaches [54, 15, 21, 29, 55] develop skills online to improve coordination. However, they do not emphasize reusing these skills for unseen tasks. ODIS [60] brings skill discovery to offline MARL, extracting generalizable skills from offline multi-task data. However, the skills discovered by ODIS are limited to the dataset they are derived from and often perform poorly on unseen tasks when the dataset quality is only moderate. Discovering high-quality reusable skills remains a significant challenge.
Hybrid Reinforcement Learning
Hybrid RL [45] has been popular recently since it can take advantage of both purely online and offline methods. Recent efforts have focused on developing offline-to-online RL, a promising paradigm to reuse offline discovered skills [12] or offline learned behaviors [19, 61]. Other studies [26, 45, 32] have concentrated on adapting Q-learning to hybrid settings. Notably, research [26, 2] on integrating offline data and online interactions into a hybrid buffer aligns closely with our approach. The work in [26] introduces a balanced replay scheme that effectively utilizes online samples by leveraging relevant, near-on-policy offline samples. RLPD [2] employs symmetric sampling, where each batch comprises 50% data from the online replay buffer and 50% from the offline data buffer. However, the application of hybrid settings to multi-agent environments is still relatively unexplored.
Background
Recent multi-task MARL works consider policy learning among two or several cooperative multi-agent tasks. In our settings, we focus on a multi-agent task set $\left\{\boldsymbol{\mathcal{T}}\right\}$ which contains tasks with varying team sizes. A multi-agent task $\mathcal{T}_{i}\in\left\{\boldsymbol{\mathcal{T}}\right\}$ can be described as a decentralized partially observable Markov decision process (Dec-POMDP) [31] consisting of a tuple $G=\left\langle I,S,A,P,\Omega,O,R,\gamma\right\rangle$. $i\in I\equiv\{1,\ldots,n\}$ is one of the agents and $s\in S$ describes the global state of the environment. At each time step, each agent $i\in I$ chooses an action $a^{i}\in A$, forming a joint action $\mathbf{a}\in\mathbf{A}\equiv A^{n}$. This causes a transition on the environment according to the state transition function $P\left(s^{\prime}\mid s,\mathbf{a}\right):S\times\mathbf{A}\times S\rightarrow[0,1]$. All agents would receive the reward according to the reward function $r(s,\mathbf{a}):S\times\mathbf{A}\rightarrow\mathbb{R}$ and $\gamma\in[0,1)$ is a discount factor. In a partially observable scenario, the agents could not get the global state $s$ but draw their individual observations $o$ $\in\Omega$ according to observation function $O(s,i):S\times I\rightarrow\Omega$. $\tau_{i}\in\boldsymbol{\tau}$ denotes the trajectory of agent $i$ which is an action-observation history $\left(o_{i}^{1},a_{i}^{1},\ldots,o_{i}^{t-1},a_{i}^{t-1},o_{i}^{t}\right)$.
We partition the multi-agent task set $\{\boldsymbol{\mathcal{T}}\}$ into source $\{\boldsymbol{\mathcal{T}}_{source}\}$ and target $\{\boldsymbol{\mathcal{T}}_{target}\}$ subsets. For source tasks, we combine their respective online interaction environments $G_{source}$ with offline datasets $\mathcal{D}$, comprising pre-collected agent trajectories $\boldsymbol{\tau}=(s,\mathbf{o},\mathbf{a},r,s^{\prime},\mathbf{o^{\prime}})$ to train our general agents. Post-training, these agents are directly deployed to execute target tasks in a zero-shot setting, without additional training or fine-tuning. Given the uncertain sources and variable quality of the offline datasets $\mathcal{D}$, they cannot be considered expert experiences.
Hybrid Training for Enhanced Multi-Task Generalization
In this section, we detail HyGen designed to enhance multi-task generalization through hybrid training. The algorithm is structured into two main components: 1) unsupervised discovery of general skills from multi-task offline datasets $\mathcal{D}$ and 2) hybrid high-level policy learning to refine and sequence the discovered skills.
Unsupervised Offline General Skill Discovery
Good skills are expected to be general latent knowledge across different tasks in multi-agent reinforcement learning. We assume the skill $z_{i}$ for agent $i$ is a discrete variable from a finite skill set $\mathcal{Z}$, where the number of skills $|\mathcal{Z}|$ is a hyper-parameter. In this project, We use a pair of VAE-style [18] networks, which contain a multi-head attention global trajectory encoder and an action decoder to abstract the skills from multi-task offline datasets. In terms of basic functions, the global trajectory encoder $q(z_{i}|s,\mathbf{a},i)$ extracts the information containing the global state $s$ and joint action $\mathbf{a}$ among each agent in the multi-task offline trajectories datasets $\mathcal{D}$ into general skill $z_{i}$ for agent $i$. There are different lengths of state $s$ and joint action $\mathbf{a}$ across different tasks. To handle this issue we reuse the task decompose module in UPDet [20]. After the decomposing, $(s,\mathbf{a})$ in different tasks is transferred into a group of entities $\mathbf{e}=\{e_{agent}^{1},e_{agent}^{2},\dots,e_{enemy}^{1},e_{enemy}^{2},\dots\}$ with same entity length. Because each entity $e^{i}$ only contains its own relevant environment information, we use $\mathcal{N}$-head self-attention to calculate the mutual influence between the entities containing all agents and the substitute entities. For each head $i$, we first compute separate query, key, and value projections: $Q_{i}=MLP_{Q}^{i}(\mathbf{e}),K_{i}=MLP_{K}^{i}(\mathbf{e}),V_{i}=MLP_{V}^{i}(\mathbf{e})$, then we can calculate the separate attention
$Attn_{i}(\mathbf{e})=\operatorname{softmax}\left(\frac{Q_{i}K_{i}^{T}}{\sqrt{d_{K_{i}}}}\right)V_{i},\quad d_{K_{i}}=\operatorname{dim}(K_{i})$
The final attention of whole entities is $Attn_{total}=\mathrm{concat}\{Attn_{1},Attn_{2},\dots,Attn_{n}\},n=\mathcal{N}$. Since self-attention computing does not change the relative position of entities in the group, we can extract the attention embedding for each agent to compute the general skill $z_{i}=MLP_{e}(Attn^{agent_{i}}),Attn^{agent_{i}}\in Attn_{total}$. It is noted that since each head in self-attention can learn to focus on different features of the input data and capture information from different representational subspaces [49], we can regard the latent information abstracted by each head as one skill. Therefore in this project, we set the number of skills $|\mathcal{Z}|$ equal to the number of attention heads $\mathcal{N}$.
Figure 2: Training framework during the general skill discovery phase of HyGen. The global trajectory encoder extracts a set of general skills common across different tasks from multi-task offline datasets, while the action decoders learn to delineate different agent actions within the discovered skills. The global trajectory encoder uses a task decomposer and multi-head self-attention to handle varying input from different tasks.
After the global trajectory encoder outputs skills, we use an action decoder to convert the skill obtained for each agent into corresponding task-specific actions. Since task-specific actions are executed in decentralized situations, acquiring global information and backward trajectory is impractical for individual agents. Therefore, the action decoder predicts a task-specific action $\hat{a}_{i}\sim p\left(\cdot\mid\tau_{i},z_{i}\right)$ using an agent $i$’s local information $\tau_{i}$ and the chosen skill $z_{i}$ output by the encoder.
Following $\beta$-VAE [18], The training objective is to maximize the likelihood of the real action $a_{i}$ from data, along with the KL divergence [17] between $q(z_{i}|s,\mathbf{a},i)$ and a uniform prior $\tilde{p}\left(z_{i}\right)$ as a regularization. The regularization with a uniform distribution of cooperative strategies can prevent the state encoder from choosing similar skills for all inputs, thereby helping to discover distinguished skills. The final objective reads:
$L_{\mathrm{VAE}}\left(\theta_{s},\phi\right)=-\mathbb{E}_{\boldsymbol{\tau}\sim\mathcal{D}}\Bigg{[}$ $\sum_{i=1}^{n}\mathbb{E}_{q(z_{i}|s,\mathbf{a},i)}\Big{[}\log p\left(a_{i}\mid\tau_{i},z_{i}\right)\Big{]}$
$-\beta D_{\mathrm{KL}}\Big{(}q(\cdot)\|\tilde{p}(\cdot)\Big{)}\Bigg{]}$
where $\theta_{s}$ and $\phi$ denote the parameters of the global trajectory encoder and the action decoder respectively, and $\beta$ is the regularization coefficient. There is no need for any external rewards during the training process, so the first stage is unsupervised training. Figure 2 summarizes the skill discovery processing.
Hybrid High-level Policy Learning
After discovering general skills from multi-task offline datasets, we further learn general high-level policy to use these skills with hybrid training. In this work, we use a hybrid sampling approach, fully utilizing trajectory data generated by online environment exploration and existing trajectories in offline datasets during the data sampling process. Contrasting with [2, 12, 19, 50], who either entirely disregard offline data or blend offline and online data uniformly, our approach dynamically adjusts the proportion of data used during training. Given that the model parameters are initially near-random, leveraging offline trajectories predominantly at the outset provides a stable starting point for learning. As the model’s performance improves, it increasingly benefits from exploring the online environment, thereby gradually transitioning to a higher proportion of online data to refine policies. When the model performance is close to the performance bottleneck of offline trajectories, the model mainly relies on exploration in the online environment to obtain better policies. Online exploration trajectories account for the majority of the training batch, and offline trajectories maintain a small proportion. In practice, We implement a linear decay scheme for adjusting the hybrid ratio $R_{h}$, defined by the following equation:
$R_{h}=\max(R_{end},\frac{(R_{start}-R_{end})\cdot t}{N})$
where $R_{start}$ is the initial hybrid ratio value, $R_{end}$ is the final and minimum hybrid ratio value, $N$ is the total number of time steps over which the hybrid ratio will decrease, and $t$ is the current time step. Each training batch comprises $R_{h}\times\mathcal{B}$ offline trajectories from dataset $\mathcal{D}$ and $(1-R_{h})\times\mathcal{B}$ online trajectories from replay buffer $\mathcal{R}$, where $\mathcal{B}$ is the batch size.
Our approach utilizes a QMIX-style value-based MARL method, as delineated by Rashid et al. [37], integrated within the Centralized Training with Decentralized Execution (CTDE) paradigm [33] to train the high-level policy. Similar to QMIX, it tries to learn a global value function $Q_{\mathrm{tot}}(\boldsymbol{\tau},\boldsymbol{z})$ that can be decomposed into agents’ individual value functions $Q_{1}\left(\tau_{1},z_{1}\right),\ldots,Q_{n}\left(\tau_{n},z_{n}\right)$. This global value function $Q_{\mathrm{tot}}(\boldsymbol{\tau},\boldsymbol{z})$ can be trained with the squared TD loss as follows:
$L_{\mathrm{TD}}\left(\theta_{v}\right)=\mathbb{E}_{\left(\boldsymbol{\tau},\boldsymbol{\tau}^{\prime}\right)\sim\mathcal{D,R}}\Bigg{[}$ $\Big{(}r+\gamma\max_{\boldsymbol{z}^{\prime}}Q_{\mathrm{tot}}\left(\boldsymbol{\tau}^{\prime},\boldsymbol{z}^{\prime};\theta_{v}^{-}\right)$
$-Q_{\mathrm{tot}}\left(\boldsymbol{\tau},\boldsymbol{z};\theta_{v}\right)\Big{)}^{2}\Bigg{]}$
Following previous MARL methods [37, 20], we use $\theta_{v}$ to denote all parameters in the value networks and the action decoder, $\theta_{v}^{-}$ to denote parameters of target networks. To address potential performance bottlenecks from sub-optimal offline data in the previous skill discovery phase, which could impair cross-task performance, we incorporate the action decoder in the hybrid training stage, enhancing skill application. When estimating Q-targets, we choose the joint general skills $z^{\prime}$ by selecting each skill $z^{\prime}_{i}$ with maximal individual Q-value $Q_{i}\left(\tau^{\prime}_{i},z^{\prime}_{i}\right)$ to avoid search in the large joint skills space, as the same as [46, 37]. Finally, We adopt a mixing network to ensure that it can satisfy the individual-global-max (IGM) [51] principle which promises the action selection with individual value functions is correct.
Figure 3: Training framework during the high-level policy learning phase of HyGen. The hybrid buffer contains trajectories from online buffer $\mathcal{R}$ and offline dataset $\mathcal{D}$. The observation encoder extracts representations from local information. Meanwhile, the mixing network employs self-attention to accommodate varying input dimensions across different tasks.
One challenge that remains is that we cannot directly get skills information since there are only state and joint actions recorded in both offline datasets and online replay buffers. Reusing skills calculated by the pre-trained global trajectory encoder are obtained with global information, which does not follow CTED. Consequently, we train a local observation encoder $\hat{q}(\cdot\mid\tau_{f}^{i})$, leveraging only agent $i$’s local trajectory, comprising its specific action sequences and local observations, to infer skills. The output distribution is expected to be similar to the pre-trained global trajectory encoder $q(z_{i}|s,\mathbf{a},i)$. We calculate the KL-divergence [17] between them to update the local observation encoder as the consistent loss $L_{\mathrm{c}}$ below:
$L_{\mathrm{c}}(\phi_{o})=\sum_{i=1}^{n}\mathbb{E}_{\boldsymbol{\tau}\sim\mathcal{D,R}}[D_{KL}(\hat{q}(\cdot\mid\tau_{f}^{i})\left|\right|q(z_{i}|s,\mathbf{a},i))]$
where $\phi_{o}$ denotes parameters of the local observation encoder in the individual value network.
Task Expert Medium
BC-best UPDeT-I UPDeT-m ODIS HyGen (ours) BC-best UPDeT-I UPDeT-m ODIS HyGen (ours)
Source Tasks
3m 96.9 $\pm$ 4 71.9 $\pm$ 14 82.1 $\pm$ 10 97.4 $\pm$ 2 99.1 $\pm$ 1 65.4 $\pm$ 14 56.6 $\pm$ 14 51.2 $\pm$ 3 85.9 $\pm$ 10 91.5 $\pm$ 11
5m6m 50.4 $\pm$ 2 12.1 $\pm$ 12 17.2 $\pm$ 28 53.9 $\pm$ 5 61.2 $\pm$ 8 21.9 $\pm$ 3 5.6 $\pm$ 4 6.3 $\pm$ 4 22.7 $\pm$ 7 31.6 $\pm$ 7
9m10m 95.3 $\pm$ 1 26.6 $\pm$ 12 3.1 $\pm$ 5 80.4 $\pm$ 8 96.4 $\pm$ 3 63.8 $\pm$ 10 34.4 $\pm$ 13 28.5 $\pm$ 10 78.1 $\pm$ 3 79.2 $\pm$ 4
Unseen Tasks
4m 92.1 $\pm$ 3 28.6 $\pm$ 21 33.0 $\pm$ 27 95.3 $\pm$ 3 95.8 $\pm$ 4 48.8 $\pm$ 21 21.6 $\pm$ 17 14.1 $\pm$ 5 61.7 $\pm$ 17 91.4 $\pm$ 8
5m 87.1 $\pm$ 10 40.1 $\pm$ 25 33.6 $\pm$ 40 89.1 $\pm$ 10 99.5 $\pm$ 1 76.6 $\pm$ 14 77.4 $\pm$ 16 67.2 $\pm$ 21 85.9 $\pm$ 11 96.5 $\pm$ 6
10m 90.5 $\pm$ 3 33.9 $\pm$ 25 54.7 $\pm$ 44 93.8 $\pm$ 2 93.5 $\pm$ 5 56.2 $\pm$ 20 36.8 $\pm$ 20 32.9 $\pm$ 11 61.3 $\pm$ 11 96.4 $\pm$ 3
12m 70.8 $\pm$ 15 10.9 $\pm$ 18 17.2 $\pm$ 28 58.6 $\pm$ 11 85.2 $\pm$ 6 24.0 $\pm$ 10 4.0 $\pm$ 5 3.2 $\pm$ 3 35.9 $\pm$ 8 81.5 $\pm$ 14
7m8m 18.8 $\pm$ 3 0.8 $\pm$ 1 0.0 $\pm$ 0 25.0 $\pm$ 15 28.9 $\pm$ 12 1.6 $\pm$ 1 2.4 $\pm$ 2 0.0 $\pm$ 0 28.1 $\pm$ 22 24.5 $\pm$ 9
8m9m 15.8 $\pm$ 3 1.6 $\pm$ 1 0.0 $\pm$ 0 19.6 $\pm$ 6 25.7 $\pm$ 9 3.1 $\pm$ 3 3.1 $\pm$ 3 2.3 $\pm$ 2 4.7 $\pm$ 2 22.3 $\pm$ 10
10m11m 45.3 $\pm$ 11 0.8 $\pm$ 1 0.0 $\pm$ 0 42.2 $\pm$ 7 57.2 $\pm$ 13 19.7 $\pm$ 8 2.4 $\pm$ 1 4.0 $\pm$ 3 29.7 $\pm$ 15 47.2 $\pm$ 13
10m12m 1.0 $\pm$ 1 0.0 $\pm$ 0 0.0 $\pm$ 0 1.6 $\pm$ 1 13.8 $\pm$ 4 0.0 $\pm$ 0 0.0 $\pm$ 0 0.0 $\pm$ 0 1.6 $\pm$ 1 5.2 $\pm$ 2
13m15m 0.0 $\pm$ 0 0.0 $\pm$ 0 0.0 $\pm$ 0 2.3 $\pm$ 2 9.5 $\pm$ 5 0.6 $\pm$ 1 0.0 $\pm$ 0 0.0 $\pm$ 0 1.6 $\pm$ 1 9.3 $\pm$ 6
Table 1: Average test win rates of the final policies in the task set marine-hard with different data qualities, averaged over five random seeds. We use abbreviated names for asymmetric tasks for simplicity. For instance, ’9m10m’ refers to the SMAC map ’9m_vs_10m’. Results of BC-best represent the highest test win rates between BC-t and BC-r.
The out-of-distribution (OOD) problem refers to the challenge of dealing with situations or state-action pairs that were not encountered in the pre-collected dataset on which the agent is trained and it directly impacts the reliability and generalization of the trained models to new, unseen environments. To tackle the out-of-distribution issue, we adopt the popular conservative Q-learning (CQL) [25] method. Different from purely offline RL, in this project, the use of offline data changes according to $R_{h}$. Therefore, when using CQL loss, $R_{h}$ is used as a coefficient to control the impact of CQL on the learning process. To be concise, the total loss term in the high-level policy learning phase is presented as
$L_{\mathrm{total}}(\theta_{v},\phi_{o})=L_{\mathrm{TD}}\left(\theta_{v}\right)+\alpha\cdot L_{\mathrm{c}}(\phi_{o})+\eta\cdot R_{h}\cdot L_{\mathrm{CQL}}$
where $\alpha$ and $\eta$ are two coefficients.
Zero-shot Execution
In zero-shot decentralized executions for test tasks, local information is employed to compute Q-values for each skill through individual value networks $Q_{i}\left(\tau_{i},z_{i}\right)$, with the optimal skill being selected based on the highest Q-value. The action decoder then utilizes this skill in conjunction with agents’ local trajectories to formulate actions tailored to the specific task, enabling effective zero-shot execution.
Experiments
In this section, we assess HyGen’s multi-task generalization capabilities, specifically focusing on zero-shot generalization across unseen tasks. Our experiments utilize custom-designed task sets from the StarCraft Multi-Agent Challenge (SMAC) [39], where we employ offline data of varied quality integrated with corresponding online interaction environments. We benchmark HyGen against purely online and offline methods across multiple source tasks, further examining its transfer capabilities in multi-scenario zero-shot transfer tasks. Experimental outcomes demonstrate that HyGen significantly outperforms purely online and offline methods.
Baselines
We compare HyGen with state-of-the-art multi-task MARL methods. Given the scarcity of such methods, we include baselines developed by ODIS [60]:
  • BC-t [48, 60], a transformer-based behavior cloning method sharing the same structure as ODIS [60].
  • BC-r [48, 60], a transformer-based behavior cloning method that incorporates return-to-go information [6] in addition to the features of BC-t.
  • UPDeT-m [20, 60], a transformer-based universal MARL model using the transformer-based mixing network of ODIS [60] to facilitate simultaneous multi-task learning.
  • UPDeT-l [20, 60], a transformer-based universal MARL model that utilizes the linear decomposable network from VDN [46] for multi-task learning.
  • ODIS [60], an offline multi-task MARL method capable of discovering general skills and learning generalizable policies, thus enabling zero-shot generalization to unseen tasks.
  • QMIX [37], a prevalent online MARL baseline operating under the CTDE paradigm, lacks zero-shot generalization capabilities. However, it still can be utilized to validate the efficiency of HyGen.
For our experiments in SMAC, we use the marine-hard and stalker-zealot task sets and both expert and medium offline datasets as defined and collected by ODIS [60], to ensure comparable fairness. Those task sets include three source tasks for training and multiple unseen tasks for evaluation. Agents are required to control various units, such as marines and stalkers, with the number of controllable agents and target enemies differing across tasks. Detailed descriptions of the task sets and properties of the offline datasets are available in Appendix A.
Experimental Results
We conduct experiments using the task sets with two different data qualities, expert and medium. We train online, offline, and our hybrid methods with offline data only from three source tasks and online environments respectively and evaluate them in a wide range of unseen tasks. To ensure comparable fairness, we train each method with the same 35k training step number. Detailed hyper-parameters of the experiments are available in Appendix B. The average test win rates of the marine-hard task set are shown in Table 1. The tables record the best test win rates between BC-t and BC-r as BC-best. We find that HyGen consistently outperforms other baselines in both source tasks and unseen tasks. HyGen can discover general skills from multi-task data and reuse them with high-level policies, resulting in superior and stable performance compared with UPDeT-l and UPDeT-m, which cannot generalize well among different levels of tasks. We find that BC methods and ODIS sometimes present comparable performance to HyGen, particularly with expert datasets. However, in real-world scenarios where non-expert data quality is more common, these purely offline methods are hampered by data quality limitations. This results in less robust performance and weaker cross-task generalization compared to HyGen, as clearly demonstrated by the training outcomes on the medium dataset. The results of the experiment for the stalker-zealot task set can be found in Appendix C.
Figure 4: Comparison of HyGen, QMIX, and ODIS on the source task $3m$. Both HyGen and ODIS were trained using medium-quality offline datasets. We run HyGen and ODIS for 35k steps due to early convergence and train QMIX for 100k steps. All experiment results were conducted over five random seeds.
We also compare the efficiency of HyGen with the online MARL method QMIX and the offline MARL method ODIS. As illustrated in Figure 4, HyGen’s learning speed surpasses that of QMIX, demonstrating greater sample efficiency. Initially, HyGen and ODIS exhibit comparable learning speeds, but after 8k steps, HyGen becomes more efficient than ODIS. We believe that this improvement can be attributed to the increasing significance of online interactions, which continually enhance performance over time, unlike purely offline MARL methods that eventually encounter dataset limitations.
Ablation Study
In our ablation studies, we investigate the effectiveness of components in our proposed HyGen structure. First, we try to find whether the linearly decreasing hybrid ratio scheme can yield better performance than the fixed. We perform HyGen hybrid training separately with the dynamic hybrid ratio and three fixed hybrid ratios which are 20%, 50%, and 80% in the marine-hard task set with medium-quality offline datasets. As we see in Figure 5, HyGen with a linearly decreasing hybrid ratio outperforms those with fixed hybrid ratios in both effectiveness and efficiency. This improvement indicates that a linearly decreasing hybrid ratio better utilizes the initial efficiency of offline learning, as offline data typically contain more useful experiences than early-stage online interactions. Starting with a higher percentage of samples from offline datasets enhances sample efficiency. Furthermore, a linearly decreasing ratio gradually increases the proportion of samples from the online replay buffer over time, progressively integrating the diversity of online interactions for exploration.
(a) Source Task: $3m$
(b) Unseen Task: $10m$
Figure 5: Average test win rates of HyGen using a linearly decreasing hybrid ratio and three fixed hybrid ratios—20%, 50%, and 80%—in the marine-hard task set with the medium dataset. All experiment results were conducted over five random seeds.
We also conduct experiments to investigate our proposed action decoder refinement during the hybrid training phase. We run experiments with two variants of HyGen, with and without action decoder refinement, during the hybrid training phase in the marine-hard task set using medium-quality offline datasets and present average test win rates in source tasks and unseen tasks. Figure 6 shows that HyGen with action decoder refinement performs better, suggesting that refining skills during hybrid training reduces their dependency on the quality of the offline dataset. This improvement occurs because online interactions contribute abundant experiences that foster exploration and enhance skill diversity.
(a) Source Task: $3m$
(b) Unseen Task: $10m$
Figure 6: Average test win rates of HyGen with or without action decoder refinement in the marine-hard task set with the medium dataset. All experiment results were conducted over five random seeds.
We then evaluate the effectiveness of the dynamic CQL loss scheme during the hybrid training phase. We conduct experiments with HyGen using the dynamic CQL loss scheme, comparing its performance against both the fixed CQL loss and no CQL loss schemes in the marine-hard task set with medium-quality offline datasets. The superior results of HyGen with the dynamic CQL loss indicate that this approach not only mitigates the OOD problem but also avoids the impact of excessive Q-value penalties on training performance in hybrid training, as shown in Figure 7.
(a) Source Task: $3m$
(b) Unseen Task: $10m$
Figure 7: Average test win rates of HyGen using dynamic, fixed, and no CQL loss scheme in the marine-hard task set with the medium dataset. All experiment results were conducted over five random seeds.
The number of general skills is a key hyper-parameter of HyGen which we recommend to set equal to the number of attention heads according to the self-attention mechanism. To substantiate this, we conducted experiments comparing performance metrics using a fixed number of attention heads with varying skill numbers in the marine-hard task set with medium-quality offline datasets. Results indicate that equalizing skill and attention head numbers yield optimal performance, as detailed in Appendix C.
Conclusion
We introduce HyGen, a novel hybrid MARL framework, Hybrid Training for Enhanced Multi-Task Generalization, which integrates online and offline learning to ensure both multi-task generalization and training efficiency. By utilizing datasets of limited offline experiences and engaging in small-scale interactive environments, HyGen effectively discovers general skills. This approach enables the learning of a general policy applicable across diverse tasks, leading to enhanced performance in both familiar source tasks and novel, unseen tasks. Our experimental results demonstrate that HyGen effectively addresses the performance limitations inherent in offline MARL algorithms and significantly outperforms traditional online MARL algorithms in terms of efficiency. We believe that the success of HyGen underscores the importance of integrating skill discovery with hybrid training methodologies to achieve generalization in cooperative MARL scenarios and will be instrumental for the practical application of MARL in real-world settings. Future work will focus on exploring integrating large language models (LLM) to utilize their universal knowledge to further enhance the efficiency and adaptability of HyGen, particularly focusing on scalability across even more diverse and dynamic real-world applications.
Appendix A A. Descriptions of Task Sets and Offline Multi-task Datasets
Task Sets
The StarCraft Multi-Agent Challenge (SMAC) [39] represents a widely recognized cooperative multi-agent testbed featuring diverse StarCraft micromanagement scenarios. This study utilizes two distinct SMAC task sets—marine-hard and stalker-zealot—each involving different agent types, defined by ODIS [60]. The marine-hard task set comprises various marine battle scenarios, wherein groups of allied marines confront equivalent or superior numbers of built-in-AI enemy marines. Conversely, the stalker-zealot task set features symmetric battles involving equal numbers of built-in-AI stalkers and zealots on opposing sides. Aiming for generalization to unseen tasks with limited offline data and online interaction environments, we designate three tasks from each set for training purposes, reserving the remainder for evaluation. Detailed attributes of these task sets are enumerated in Table 2 and Table 3.
Task type Task Ally units Enemy units Properties
Source tasks 3 m 3 Marines 3 Marines homogeneous & symmetric
5m_vs_6m 5 Marines 6 Marines homogeneous & asymmetric
9m_vs_10m 9 Marines 10 Marines homogeneous & asymmetric
Unseen tasks 4 m 4 Marines 4 Marines homogeneous & symmetric
5 m 5 Marines 5 Marines homogeneous & symmetric
10 m 10 Marines 10 Marines homogeneous & symmetric
12 m 12 Marines 12 Marines homogeneous & symmetric
7m_vs_8m 7 Marines 8 Marines homogeneous & asymmetric
8m_vs_9m 8 Marines 9 Marines homogeneous & asymmetric
10m_vs_11m 10 Marines 11 Marines homogeneous & asymmetric
10m_vs_12m 10 Marines 12 Marines homogeneous & asymmetric
13m_vs_15m 13 Marines 15 Marines homogeneous & asymmetric
Table 2: Descriptions of tasks in the marine-hard task set. [60]
Task type Task Ally units Enemy units Properties
Source tasks 2s3z
2 Stalkers,
3 Zealots
2 Stalkers,
3 Zealots
heterogeneous & symmetric
2s4z
2 Stalkers,
4 Zealots
2 Stalkers,
4 Zealots
heterogeneous & symmetric
3s5z
3 Stalkers,
5 Zealots
3 Stalkers,
5 Zealots
heterogeneous & symmetric
Unseen tasks 1s3z
1 Stalkers,
3 Zealots
1 Stalkers,
3 Zealots
heterogeneous & symmetric
1s4z
1 Stalkers,
4 Zealots
1 Stalkers,
4 Zealots
heterogeneous & symmetric
1s5z
1 Stalkers,
5 Zealots
1 Stalkers,
5 Zealots
heterogeneous & symmetric
2s5z
2 Stalkers,
5 Zealots
2 Stalkers,
5 Zealots
heterogeneous & symmetric
3s3z
3 Stalkers,
3 Zealots
3 Stalkers,
3 Zealots
heterogeneous & symmetric
3s4z
3 Stalkers,
4 Zealots
3 Stalkers,
4 Zealots
heterogeneous & symmetric
4s3z
4 Stalkers,
3 Zealots
4 Stalkers,
3 Zealots
heterogeneous & symmetric
4s4z
4 Stalkers,
4 Zealots
4 Stalkers,
4 Zealots
heterogeneous & symmetric
4s5z
4 Stalkers,
5 Zealots
4 Stalkers,
5 Zealots
heterogeneous & symmetric
Table 3: Descriptions of tasks in the stalker-zealot task set. [60]
Offline Multi-task Datasets
As stated in the experiments section, we utilize the same offline dataset as ODIS [60] to maintain fairness in our evaluations. Definitions of expert and medium qualities are listed below:
  • The expert dataset contains trajectory data collected by a QMIX policy trained with 2, 000, 000 steps of environment interactions. The test win rate of the trained QMIX policy (as the expert policy) is recorded for constructing medium datasets.
  • The medium dataset contains trajectory data collected by a QMIX policy (as the medium policy) whose test win rate is half of the expert QMIX policy.
Considering our focus on generalizing to unseen tasks, we employ offline datasets exclusively from the source tasks in the three aforementioned task sets. The Properties of offline datasets with different qualities are detailed in Table 4. Data from various tasks are amalgamated into a multi-task dataset, facilitating simultaneous multi-task policy training.
Task Quality Trajectories Average win rate (%) Average return
3m expert 2000 99.10 19.8929
medium 2000 54.02 13.9869
5m_vs_6m expert 2000 71.85 17.3424
medium 2000 27.51 12.6408
9m_vs_10m expert 2000 94.31 19.6140
medium 2000 41.46 15.5049
2s3z expert 2000 96.02 19.7655
medium 2000 44.65 16.6279
2s4z expert 2000 95.09 19.7402
medium 2000 49.65 16.8735
3s5z expert 2000 95.18 19.7850
medium 2000 31.14 16.3126
Table 4: Properties of offline datasets with different qualities. [60]
Appendix B B. Experiments Details
The specific hyper-parameters of HyGen are listed in Table 5. All the tabular results show the performance of HyGen with 50, 000 optimization steps, and the steps of the hybrid high-level policy learning phase are the subtraction of the general skill discovery steps from the total steps. Our experiments are conducted on a server equipped with one Intel Xeon E5 CPU@3.60GHz processor (6 cores, 12 threads), 128 GB memory, and 2 RTX 3090 GPU cards and it usually costs 10-14 hours typically. Our HyGen code follows Apache License 2.0, the same as the PyMARL framework.
Hyper-parameters Value
hidden layer dimension 64
attention embedding length 128
$\alpha$ 5.0
$\beta$ 0.001
$\eta$ 5.0
$\|\mathcal{Z}\|$ 4
number of attention heads $\mathcal{N}$ 4
steps of general skill discovery 15000
steps of high-level policy learning 35000
$R_{start}$ 1.0
$R_{end}$ 0.1
linear decay steps 5000
batch size $\mathcal{B}$ 32
optimizer Adam
learning rate 0.0005
Table 5: Hyper-parameters of HyGen.
Appendix C C. Additional Experiments
Experiments with different skill numbers
The number of general skills, a critical hyper-parameter in HyGen, is recommended to be set equal to the number of attention heads to align with the self-attention mechanism’s design. To validate this approach, we conducted experiments within the marine-hard task set, utilizing medium-quality offline datasets and comparing performance across a fixed number of attention heads with varying numbers of general skills. Table 6 displays the average test win rates for policies trained with various general skills counts, each within a configuration of four fixed attention heads and medium data quality. Results indicate that a general skill count of four yields comparable performances across most unseen tasks, suggesting that HyGen can effectively abstract latent information from each attention head into general skills. Conversely, a general skill count that is either too low or too high compromises generalization to unseen tasks, due to either an overload or a deficit of information encapsulated within each skill.
Task skill num.1 skill num.2 skill num.3 skill num.4 skill num.5 skill num.6 skill num.7 skill num.8
Source Tasks
3m 60.4 $\pm$ 7 74.6 $\pm$ 11 81.5 $\pm$ 16 91.5 $\pm$ 11 85.4 $\pm$ 4 86.6 $\pm$ 11 91.2 $\pm$ 3 85.9 $\pm$ 10
5m6m 19.8 $\pm$ 2 21.1 $\pm$ 12 22.2 $\pm$ 8 31.6 $\pm$ 7 30.9 $\pm$ 3 25.6 $\pm$ 14 27.5 $\pm$ 6 20.7 $\pm$ 4
9m10m 58.4 $\pm$ 6 56.6 $\pm$ 12 73.1 $\pm$ 5 79.2 $\pm$ 4 73.8 $\pm$ 10 74.4 $\pm$ 13 64.9 $\pm$ 9 72.5 $\pm$ 13
Unseen Tasks
4m 46.8 $\pm$ 3 78.6 $\pm$ 12 73.0 $\pm$ 17 91.4 $\pm$ 8 88.8 $\pm$ 11 81.6 $\pm$ 7 81.9 $\pm$ 14 73.1 $\pm$ 4
5m 64.7 $\pm$ 10 80.1 $\pm$ 9 90.3 $\pm$ 9 96.5 $\pm$ 6 96.6 $\pm$ 4 87.4 $\pm$ 6 81.2 $\pm$ 11 83.0 $\pm$ 11
10m 61.4 $\pm$ 3 63.9 $\pm$ 25 84.7 $\pm$ 10 96.4 $\pm$ 3 96.2 $\pm$ 6 92.1 $\pm$ 7 93.4 $\pm$ 4 91.0 $\pm$ 5
12m 44.6 $\pm$ 12 60.9 $\pm$ 8 77.2 $\pm$ 8 81.5 $\pm$ 14 74.0 $\pm$ 10 74.4 $\pm$ 5 72.7 $\pm$ 7 66.9 $\pm$ 7
7m8m 8.8 $\pm$ 3 10.8 $\pm$ 3 21.2 $\pm$ 6 24.5 $\pm$ 9 18.6 $\pm$ 6 22.4 $\pm$ 4 12.4 $\pm$ 8 14.6 $\pm$ 3
8m9m 5.8 $\pm$ 3 11.6 $\pm$ 6 19.0 $\pm$ 6 22.3 $\pm$ 10 19.1 $\pm$ 8 23.1 $\pm$ 13 17.7 $\pm$ 9 13.1 $\pm$ 11
10m11m 15.3 $\pm$ 11 38.8 $\pm$ 4 35.1 $\pm$ 11 47.2 $\pm$ 13 29.7 $\pm$ 9 32.4 $\pm$ 14 27.3 $\pm$ 16 27.7 $\pm$ 11
10m12m 1.0 $\pm$ 1 0.0 $\pm$ 0 0.0 $\pm$ 0 5.2 $\pm$ 2 2.5 $\pm$ 5 3.7 $\pm$ 3 0.0 $\pm$ 0 0.0 $\pm$ 0
13m15m 0.0 $\pm$ 0 0.0 $\pm$ 0 0.0 $\pm$ 0 9.3 $\pm$ 6 5.6 $\pm$ 13 0.0 $\pm$ 0 0.0 $\pm$ 0 0.0 $\pm$ 0
Table 6: In the context of zero-shot execution, we assessed the average test win rates of final policies trained with varying numbers of general skills within a task set configured with 4 fixed attention heads and medium data quality. These performance evaluations are derived from averages across five random seeds. For ease of reference, asymmetric task names are abbreviated, with ’5m6m’ denoting the SMAC map ’5m_vs_6m’.
Experiments with the stalker-zealot task set
Task Expert Medium
BC-best UPDeT-I UPDeT-m ODIS HyGen (ours) BC-best UPDeT-I UPDeT-m ODIS HyGen (ours)
Source Tasks
2s3z 93.1 $\pm$ 4 53.1 $\pm$ 39 50.0 $\pm$ 33 97.7 $\pm$ 2 97.1 $\pm$ 3 48.8 $\pm$ 9 30.6 $\pm$ 12 35.0 $\pm$ 23 49.2 $\pm$ 8 73.5 $\pm$ 11
2s4z 78.1 $\pm$ 8 48.4 $\pm$ 24 23.4 $\pm$ 26 60.9 $\pm$ 6 86.2 $\pm$ 10 12.5 $\pm$ 8 28.8 $\pm$ 4 18.8 $\pm$ 10 32.8 $\pm$ 12 51.3 $\pm$ 8
3s5z 92.5 $\pm$ 4 40.6 $\pm$ 11 17.2 $\pm$ 19 87.5 $\pm$ 9 88.9 $\pm$ 13 24.4 $\pm$ 12 15.0 $\pm$ 10 25.6 $\pm$ 24 28.9 $\pm$ 6 52.6 $\pm$ 13
Unseen Tasks
1s3z 45.6 $\pm$ 23 26.6 $\pm$ 25 1.6 $\pm$ 1 76.6 $\pm$ 3 84.1 $\pm$ 5 21.9 $\pm$ 37 33.1 $\pm$ 18 3.8 $\pm$ 5 41.4 $\pm$ 18 54.2 $\pm$ 8
1s4z 60.0 $\pm$ 32 37.5 $\pm$ 31 26.6 $\pm$ 19 17.2 $\pm$ 10 44.5 $\pm$ 9 6.2 $\pm$ 7 35.0 $\pm$ 7 2.5 $\pm$ 3 50.7 $\pm$ 7 67.3 $\pm$ 7
1s5z 45.6 $\pm$ 26 14.8 $\pm$ 13 29.7 $\pm$ 26 2.5 $\pm$ 2 47.2 $\pm$ 13 3.1 $\pm$ 2 13.1 $\pm$ 11 5.0 $\pm$ 4 14.1 $\pm$ 8 34.2 $\pm$ 13
2s5z 75.6 $\pm$ 11 27.3 $\pm$ 19 23.4 $\pm$ 22 27.3 $\pm$ 6 72.4 $\pm$ 15 14.4 $\pm$ 9 17.5 $\pm$ 9 16.9 $\pm$ 14 32.0 $\pm$ 4 43.7 $\pm$ 5
3s3z 80.6 $\pm$ 9 49.2 $\pm$ 25 20.3 $\pm$ 10 89.1 $\pm$ 5 93.3 $\pm$ 6 45.6 $\pm$ 14 23.8 $\pm$ 6 24.4 $\pm$ 28 23.4 $\pm$ 9 41.3 $\pm$ 7
3s4z 92.5 $\pm$ 5 59.4 $\pm$ 16 12.5 $\pm$ 19 96.9 $\pm$ 2 93.9 $\pm$ 5 40.0 $\pm$ 19 17.5 $\pm$ 10 28.8 $\pm$ 31 50.8 $\pm$ 15 71.6 $\pm$ 9
4s3z 67.5 $\pm$ 19 50.8 $\pm$ 24 6.2 $\pm$ 4 64.1 $\pm$ 13 74.9 $\pm$ 13 28.8 $\pm$ 26 3.1 $\pm$ 4 11.2 $\pm$ 18 13.3 $\pm$ 7 52.6 $\pm$ 13
4s4z 53.1 $\pm$ 18 41.4 $\pm$ 16 7.8 $\pm$ 13 79.7 $\pm$ 10 74.1 $\pm$ 16 20.0 $\pm$ 12 1.9 $\pm$ 2 1.2 $\pm$ 1 12.5 $\pm$ 7 44.2 $\pm$ 15
4s5z 40.6 $\pm$ 19 28.1 $\pm$ 17 5.5 $\pm$ 7 86.7 $\pm$ 12 89.9 $\pm$ 7 14.4 $\pm$ 8 5.0 $\pm$ 5 5.6 $\pm$ 8 7.0 $\pm$ 4 28.1 $\pm$ 13
4s6z 48.1 $\pm$ 23 10.9 $\pm$ 7 4.7 $\pm$ 6 88.3 $\pm$ 8 86.1 $\pm$ 12 3.8 $\pm$ 3 2.5 $\pm$ 2 1.9 $\pm$ 2 1.6 $\pm$ 1 23.8 $\pm$ 14
Table 7: Average test win rates of the final policies in the task set stalker-zealot with different data qualities, averaged over five random seeds. Results of BC-best represent the highest test win rates between BC-t and BC-r. For ease of reference, asymmetric task names are abbreviated, with ’2s3z’ denoting the SMAC map ’2s_vs_3z’.
Table 7 records the average test win rates of the stalker-zealot task set. Table records the best test win rates between BC-t and BC-r as BC-best. We find that HyGen consistently outperforms other baselines in both source tasks and unseen tasks. HyGen can discover general skills from multi-task data and reuse them with high-level policies, resulting in superior and stable performance compared with UPDeT-l and UPDeT-m, which cannot generalize well among different levels of tasks. We find that BC methods and ODIS sometimes present comparable performance to HyGen, particularly with expert datasets. However, in real-world scenarios where non-expert data quality is more common, these purely offline methods are hampered by data quality limitations. This results in less robust performance and weaker cross-task generalization compared to HyGen, as clearly demonstrated by the training outcomes on the medium dataset.